# Lightweight Multiscale Feature Fusion Super-Resolution Network Based on Two-branch Convolution and Transformer

### v1

@@ Abstract
The single image super-resolution(SISR) algorithms under deep learning currently have two main models, one based on convolutional neural networks and the other based on Transformer. The former uses the stacking of convolutional layers with different convolutional kernel sizes to design the model, which enables the model to better extract the local features of the image; the latter uses the self-attention mechanism to design the model, which allows the model to establish long-distance dependencies between image pixel points through the self-attention mechanism and then better extract the global features of the image. However, both of the above methods face their problems. Based on this, this paper proposes a new lightweight multi-scale feature fusion network model based on two-way complementary convolutional and Transformer, which integrates the respective features of Transformer and convolutional neural networks through a two-branch network architecture, to realize the mutual fusion of global and local information. Meanwhile, considering the partial loss of information caused by the low-pixel images trained by the deep neural network, this paper designs a modular connection method of multi-stage feature supplementation to fuse the feature maps extracted from the shallow stage of the model with those extracted from the deep stage of the model, to minimize the loss of the information in the feature images that is beneficial to the image restoration as much as possible, to facilitate the obtaining of a higher-quality restored image. The practical results finally show that the model proposed in this paper is optimal in ×$\times 2$, ×$\times 3$ and ×$\times 4$ image recovery performance when compared with other lightweight models with the same amount of parameters.
@@ I Introduction
Figure 1: Visual comparison of other models with our model in terms of number of parameters and PSNR values
@@
In addition, to improve the model’s ability to learn image features in the training phase and to reduce the computational effort in the inference phase, this paper adopts the design idea of repVGG[11] on the convolutional branch. The model makes the convolutional branch have three branches in each module in the middle by adding another $1\times 1$ convolutional branch in addition to the residual link. This design increases the complexity of the model and can improve the model’s learning of image features. In the inference stage of the model, the model then removes the redundant $1\times 1$ convolutional branches, which allows the model to reduce the amount of computation in inference significantly and maintain a relatively stable image recovery performance.
@@
II Related Works
II-A Lightweight Single Image Super-Resolution Algorithms
Dong et al[7]. proposed for the first time a single-image super-resolution model based on convolutional neural networks for low-resolution image recovery, called SRCNN (Super-Resolution Using Convolutional Neural Network). This model uses deep learning methods to process the recovery of low-resolution images for the first time and has achieved great success. After this, a large number of methods using deep learning and convolutional neural networks to deal with single-image super-resolution problems have appeared at home and abroad. For a general convolutional neural network model, as the number of layers and parameters increases, the expression ability and performance of the model get better. However, this way of improving model performance has its limitations because as the network size increases, a large number of parameters and computation costs will limit the deployment and training of the model on hardware devices. Therefore, to solve this problem, many lightweight approaches based on convolutional neural networks have emerged in the field of single-image super-resolution, including but not limited to the use of recursive structures[12], group convolution[13], etc. to reduce the number of parameters and computation of the model. In this paper, some of the above methods are also used to reduce the number of parameters and computation of convolutional neural networks by using depth-separable convolution, which improves the performance and computational efficiency of the model.
II-B Image Degradation Model
In single-image super-resolution studies, it is usually assumed that the low-resolution image is obtained from the high-resolution image through image degradation operations. It is thus assumed that the model can easily establish a mapping relationship between low-resolution images and high-resolution images of the same content, which also helps the subsequent supervised training. Based on this, the image recovery process of a single-image super-resolution model can be regarded as the inverse process of image degradation operation and also can be regarded as the process of mining the mapping relationship between low-resolution images and high-resolution images.
Image degradation[1] usually contains the following factors: downsampling, blurring, geometric transformation and noise. The high-resolution image is degraded by selecting one or a combination of these factors to obtain the corresponding low-resolution image. The downsampling method generally uses double and triple downsampling. Blurring is divided into motion blur and optical blur. Motion blur refers to the blur caused by the motion of the imaging device (e.g. camera, mobile phone) relative to the photographed object, while optical blur refers to simple Gaussian blur. Geometric transformations refer to the translation and rotation of an image. Noise, on the other hand, is classified according to the temporal distribution as Gaussian, gamma and exponential noise and other related noises.
Formulas can provide a brief overview of mass reduction operations:
$\begin{split}y=H(x)+n=(x\otimes k)\downarrow_{s}+n\end{split}$ (1)
where $H(\cdot)$ is the degeneracy function and $n$ denotes the noise. The degradation function $H(x)$ is obtained from a high resolution image $x$ and a fuzzy kernel $k$ It is obtained by performing a convolution operation and then downsampling. The $\downarrow_{s}$ indicates that the downsampling coefficient is $s$.
Figure 5: SESAB network architecture
III Method
@@
III-A Double-Branch Feature Fusion Block Design
Figure 7: Comparison of square-window self-attention calculation and axial-window self-attention calculation mechanisms
@@ III-A Double-Branch Feature Fusion Block Design
III-A1 Axis Transformer Block
@@
III-A2 Space Enhanced Self-Attention Block
@@
III-B Feature Reuse Block
@@
Figure 12: Visual comparison of bicubic image SR ($\times$4) methods.
@@
IV Experiments Detail
IV-A Dataset
IV-A1 Training Dataset
In this paper, we use the DIV2K image set[23], which is popular in the field of image super-resolution, as a training set, and use this dataset to train and stage validate the designed model. The dataset consists of 800 training images (image number: 1-800) and 100 validation images (image number: 801-900). To facilitate quick validation during training and intermediate model saving according to the given rules, the model uses 10 of the validation images for actual validation during training (image numbers: 841-850). The model uses double-triple downsampling for the original HR images. In addition, in order to increase the complexity of the test images and the actual effect of the model after training, the experiment also used data augmentation on the training set to increase the difficulty of the model in learning the samples by randomly rotating the training set images by $90^{\circ}$, $180^{\circ}$, $270^{\circ}$, and horizontally flipping them to enhance the expressive ability of the model.
During training, the model with magnification of ×$\times 2$ is trained first, and after the ×$\times 2$ model is trained, the trained model is loaded as a pre-trained model onto the models with magnification of ×$\times 3$ and ×$\times 4$ for subsequent training.
IV-A2 Test Dataset
To ensure the objectivity of the comparison with other models, the following experiments are also validated using the five types of standard validation sets widely used in the field: Set5[24], Set14[25], B100[26], Urban100[27], and Manga109[6].To make a fair comparison with the previous models, the peak signal to noise ratio (PSNR: peak signal to noise ratio) and structural similarity index (SSIM: structural similarity index) are used as metrics for each test set, respectively, for the brightness (Y) channel in YCbCr space. signal to noise ratio (PSNR: peak signal to noise ratio) and structural similarity (SSIM: structural similarity index) as metrics, and the PSNR and SSIM values of each test set are calculated separately and compared.
IV-B Experimental Setup
During model training, 48 $64\times 64$ image blocks were randomly selected in each batch. Adam is used as the optimiser and $\beta_{1}=0.9,\beta_{2}=0.99$ is set. All models in this paper are set to 1,000,000 iteration cycles, and the learning rate is set to $42e-$ 4 at the beginning and halved at the 50,000th, 80,000th, 90,000th, and 95,000th iterations. The loss function is set to the L1 loss function.
In this paper, we use the Pytorch framework, 2 Nvidia 2080Ti GPUs for model training and testing.
IV-C Results on Image SR
IV-C1 Quantitative comparison
@@ IV-C2 Visual Comparison
IV-D Ablation Study and Discussion
In this paper, we first remove the ATB branch and use the same structure as the residual convolution branch to replace the ATB branch, and use the two paths of the residual convolution combination to perform the extraction of low-resolution image features as well as image recovery, and the results of the image recovery are shown in the second row of Table II.
TABLE II: Ablation study on Two-Branch design
scale SESAB ATB PSNR SSIM
×$\times 4$ w w 31.24 0.9170
w w/o 30.37 0.9068
w/o w 30.55 0.9083
Afterwards, the experiment removes the SESAB branch and uses the same structure as the ATB branch to replace the residual convolutional branch, and uses the two pathways combined with the ATB branch to extract low-resolution image features for subsequent image recovery, and the results of image recovery are shown in the third row of Table II. The experimental results show that the hybrid two-way network model has better results than the two-way network with all Transformer branches and the two-way network with all Convolutional Neural Network branches.
TABLE III: Ablation study on DFB design
scale DFB PSNR SSIM
×$\times 4$ w 31.24 0.9170
w/o 31.15 0.9164
The second set of experiments (Table III) examined the effect of the presence or absence of the DFB module on the effectiveness of image recovery for the network model. This experiment removes the DFB module and replaces it with the use of a convolution kernel for the convolution along the channels after splicing along the channel dimensions, compressing the feature map from 2C to C, and then proceeding to the feature extraction of the subsequent images. It can be seen that the two-way network with the DFB module can better fuse the feature information of the two branches to achieve better image super-resolution reconstruction.
TABLE IV: Ablation study on FRB design
scale FRB PSNR SSIM
×$\times 4$ w 31.24 0.9170
w/o 31.03 0.9141
The third set of experiments (Table IV) examined the effect of the presence or absence of the FRB module on the image recovery results. In the experiments, the FRB module is removed and the convolution kernel is used for the convolution along the channels after splicing along the channel dimensions, and the final feature image is compressed from 4C to C, and then the subsequent image recovery is performed. It can be seen that the network with the FRB module is more capable of combining the information contained in the shallow and deep features to achieve better image recovery.
V Conclusion

### v2

# Lightweight single-image super-resolution network based on dual paths

@@ Abstract
Currently, there are two main mainstream models for single-image super-resolution (SISR) algorithms under deep learning, one is based on convolutional neural network (CNN) and the other is based on Transformer.The former designs the model by stacking convolutional layers with different convolutional kernel sizes, so that the model can better extract the local features of the image; the latter designs the model by using the self-attention mechanism, through which the self-attention The latter uses the self-attention mechanism to design a model that establishes long-distance dependencies between image pixels, so as to better extract the global features of an image. However, both methods face their own problems. Based on this, a new lightweight multi-scale feature fusion network model, TCC-TransNet (or CT-FusionNet), is proposed in this paper, which fuses the respective features of Transformer and Convolutional Neural Networks through a two-branch network architecture to achieve global and local information of global and local information. Meanwhile, considering the partial information loss that may be caused when deep neural networks train low-pixel images, this paper designs a modular connectivity method with multi-stage feature supplementation (MCM-FS), which fuses the feature maps extracted from the shallow stage of the model with those extracted from the deeper stage, in order to minimise the loss of feature image information conducive to image restoration, thus facilitating the acquisition of a higher-quality restored image ( RIQ). The experimental results show that compared with other lightweight models with the same number of parameters, the model proposed in this paper exhibits optimality in ×2, ×3 and ×4 image recovery performance
@@ I Introduction
II Method
@@
II-A Double-Branch Feature Fusion Block Design
@@
II-A1 Axis Transformer Block
@@
II-A2 Space Enhanced Self-Attention Block
@@
II-B Feature Reuse Block
@@
II-C Results on Image SR
II-C1 Quantitative comparison
@@ II-C2 Visual Comparison
III Conclusion